# ViMoE: An Empirical Study of Designing Vision Mixture-of-Experts

### v1

@@ 3.2 Efficient Exploration Based on Stability
Routing Degree. Another interesting observation is that the number of MoE layers $L$ required varies with the number of experts $N$. We suggest this is related to the routing degree, which represents the number of possible expert combinations and can be simply defined as $D=(C^{k}_{N})^{L}$. Since we fix the gating selection to top-$1$ (i.e., $k=1$), we obtain $D=(C^{1}_{2})^{5}=32$ for $N=2$, $D=(C^{1}_{4})^{3}=64$ for $N=4$, and $D=(C^{1}_{8})^{2}=64$ for $N=8$. This implies that approximately 32 to 64 routing combinations are sufficient for effectively partitioning and processing the data. Fewer combinations may affect performance, while more do not yield further significant gains. From another perspective, if we view the gating network allocating experts to data as a clustering process, the routing degree essentially reflects the number of clusters formed from the dataset. Each expert combination can then specialize in learning from the samples of its corresponding cluster, facilitating the model in reaching optimal effectiveness. Our results validate that end-to-end training can effectively achieve this clustering effect, without the need for additional clustering strategies to provide prior information for the gating mechanism (Liu et al., 2024).
@@
Efficient ViMoE. The above conclusions are drawn from scanning the number of MoE layers. From another perspective, we can approximately predict the routing degree by observing the expert allocation in each layer. As illustrated in Fig. 5, the routing heatmap provides evidence of which MoE layers play a critical role, potentially indicating the necessary expert combinations that impact the results. These insights guide us in refining the structural design, retaining the essential MoE layers while removing the unnecessary ones, thereby developing a more efficient ViMoE. Moreover, we expect these findings are not limited to the ImageNet-1K dataset. In Sec. 4.2, we further explore the transfer of these insights to CIFAR100 (Wang et al., 2017) to validate their generality.
@@
In Table 2, we present various ViMoE configurations and compare their parameter counts. Although sparse MoE layers increase the total number of parameters, since we set the gate to route each image to the top-1 expert, it achieves higher accuracy without increasing the activated parameter counts or the inference burden. With the inclusion of the shared expert, we further improve accuracy at relatively low extra cost. For example, when $N=8$ and $L=2$, only 2.4M additional activated parameters are required to surpass the baseline by 1.1% in accuracy. Furthermore, a comparison with $L=12$ highlights the efficiency of our structural design for ViMoE, significantly reducing parameter count without sacrificing accuracy.
@@
$N$ $L$ w/ Shared Expert Total Param. Activate Param. FLOPs Acc.
- 0 - 22.0M 22.0M 5.53G 83.1
2 5 27.9M 22.0M 5.53G 83.6
2 5 33.8M 27.9M 7.04G 84.3
2 12 50.4M 36.2M 9.17G 84.2
4 3 32.7M 22.0M 5.53G 83.9
4 3 36.2M 25.6M 6.44G 84.2
4 12 78.8M 36.2M 9.17G 84.2
8 2 38.6M 22.0M 5.53G 83.9
8 2 40.9M 24.4M 6.13G 84.2
8 12 135.5M 36.2M 9.17G 84.3
Table 2: Top-1 accuracy on ImageNet-1K. All models are evaluated at resolutions $224\times 224$. We select $N=8$, $L=2$ as a representative configuration to report. indicates the inclusion of the shared expert.
Method Arch. Activate Param. FLOPs Acc.
DINO ViT-S/16 22.1M 4.25G 81.5
BEiT ViT-S/16 22.1M 4.25G 81.7
iBOT ViT-S/16 22.1M 4.25G 82.3
DINOv2 ViT-S/14 22.0M 5.53G 83.1
DINO ViT-B/16 86.6M 17.58G 82.8
MoCov3 ViT-B/16 86.6M 17.58G 83.2
BEiT ViT-B/16 86.6M 17.58G 83.4
MAE ViT-B/16 86.6M 17.58G 83.6
iBOT ViT-B/16 86.6M 17.58G 84.0
ViMoE ViT-S/14 22.0M 5.53G 83.9
ViMoE ViT-S/14 24.4M 6.13G 84.2
Table 3: Comparison between dense structure and sparse MoE. For dense structures, $L$ indicates that each of the last $L$ layers contains two FFNs.
Arch. $L$ $N$ Activate Param. FLOPs Acc.
Dense 0 - 22.0M 5.53G 83.1
Dense 2 - 24.4M 6.13G 83.6
Dense 3 - 25.6M 6.44G 83.8
Dense 5 - 27.9M 7.04G 83.8
Dense 12 - 36.2M 9.17G 83.9
Sparse 2 8 24.4M 6.13G 84.2
Sparse 3 4 25.6M 6.44G 84.2
Sparse 5 2 27.9M 7.04G 84.3
Table 4: Ablation studies of different routing strategies. We calculate the average number of routed experts and activated parameters per image, with the total number of experts being $(N+1)\times L$ (including the shared expert).
Strategy $L$ $N$ Avg. # Experts Activate Param. Acc.
Token 2 8 16.3 38.9M 84.1
Token 3 4 14.4 35.5M 84.2
Token 5 2 14.8 33.6M 84.1
Image 2 8 4 24.4M 84.2
Image 3 4 6 25.6M 84.2
Image 5 2 10 27.9M 84.3
@@ 4.1 Image Classification on ImageNet-1K
Implementation Details. All experiments are conducted on the DINOv2 (Oquab et al., 2023) pre-trained ViT-S/14 (Dosovitskiy, 2020) and fine-tuned on ImageNet-1K (Deng et al., 2009) with $224\times 224$ image resolution for $200$ epochs. By default, we use the AdamW (Sun et al., 2021) optimizer with a batch size of $1024$, a weight decay of $0.05$, and a layer-wise learning rate decay of $0.65$. The peak learning rate is set to $1e^{-4}$ with a warm-up of $20$ epochs. For the MoE layers, we configure three different numbers of experts ($N=2$, $N=4$, and $N=8$), selecting the top-1 expert, with the load balancing loss coefficient $\alpha$ set to $0.01$.
@@
Results. Most of the empirical results on the ImageNet-1K benchmark have already been presented earlier. Here, we compare ViMoE against various self-supervised models (Bao et al., 2021; Zhang et al., 2022; Zhou et al., 2021; Oquab et al., 2023; Xinlei et al., 2021; He et al., 2022). As shown in Table 2, ViMoE achieves an 83.9% top-1 accuracy, which is 0.8% higher than DINOv2 without increasing activated parameters. With shared expert, the accuracy further improves to 84.2%, outperforming DINOv2 by 1.1%. Notably, we achieve this performance using only ViT-S/14, surpassing other methods based on ViT-B/16, while activating less than one-third of their parameters.
@@
Comparison with Dense Structures. Previous results validate the advantage of the MoE structure over dense models. However, when we introduce the shared expert, the number of activated parameters increases. To ensure fairness, we attempt to modify the DINOv2 baseline by aligning the number of activated parameters while maintaining a dense architecture. One feasible approach is to mimic the MoE by setting two experts and selecting the top-2, which allows for the addition of an extra FFN in the ViT block.
In Table 4, we present the results of dense structure with different layer counts and compare them with sparse MoE. While increasing the number of parameters provides accuracy gains, the sparse structures are obviously more efficient and have a higher upper bound. For instance, at $L=2$ with 24.4M activated parameters, sparse MoE outperforms the dense one by 0.6%.
@@
Routing Distribution. In Sec. 2.1, we introduce the load balancing loss to assist in training sparse MoE models. Its purpose is to ensure that multiple experts receive inputs more evenly, preventing the majority of data from being routed to a single expert and thus avoiding the model from degrading into a dense structure. We calculate the proportion of data allocated to each expert in the MoE layers, as shown in Fig. 6. It is evident that the gating network distributes the data relatively evenly across multiple experts. Combined with the observations from Fig. 5, this validates the expectation that MoE layers enable different experts to handle specific information.
Routing Strategy. In Sec. 2.2, we introduce the routing strategy, where experts are selected for the entire image rather than for each token. In Table 4, we conduct an ablation study comparing these two strategies, showing no significant difference in accuracy. This indicates that the image-level strategy, while simpler, is effective because it aligns with the task objective of image classification. Additionally, we calculate the average number of routed experts and activated parameters per image, further confirming that our choice is more efficient.
@@ 4.2 Validation on CIFAR100
The above-mentioned observations and conclusions are based on ImageNet-1K (Deng et al., 2009). To demonstrate generalizability, we conduct validation on CIFAR100 (Wang et al., 2017) and aim to identify the most suitable ViMoE configuration.
@@
Implementation Details. All models are fine-tuned on CIFAR100 for $100$ epochs with a weight decay of $0.3$. The peak learning rate is set to $3e^{-4}$ with a warm-up of $3$ epochs, while all other settings remain consistent with those adopted on ImageNet-1K.
Baseline and Stable ViMoE. First, we use the DINOv2 (Oquab et al., 2023) self-supervised pre-trained ViT-S/14 (Dosovitskiy, 2020) and fine-tune it on CIFAR100 as the baseline, which achieves a top-1 accuracy of 91.3%. Next, we convert the ViT blocks into the ViMoE framework. Considering that CIFAR100 has fewer classes and samples than ImageNet-1K, we set the number of experts to $N=4$ in our experiments. Based on prior experience, ViMoE with the shared expert tends to yield stable results, allowing us more flexibility in setting the number of MoE layers. We opt for the most straightforward approach by adding MoE layers to every block, i.e., $L=12$. Under this setting, ViMoE achieved a top-1 accuracy of 91.6%, surpassing the baseline by 0.3%. Additionally, we compare the model without the shared expert, which yields an accuracy of only 78.4%, falling far short of the baseline. This demonstrates that MoE is not a simple design that guarantees stable gains. In fact, the optimization complexity introduced by sparse structures in certain ViT blocks may have significant negative impacts, further highlighting the necessity of designing a stable ViMoE.
Figure 7: Routing heatmap for the $l$-th MoE layer on CIFAR100. The $x$-axis is the class ID, and the $y$-axis is the expert ID. The label order in each figure is adjusted for better readability. Darker colors indicate a higher proportion of images from the corresponding class routed to the expert.
Efficient Structures Derived from Observations. We observe the behavior of MoE within the stable ViMoE and further analyze which layers play a critical role. Following the approach outlined in Sec. 3.2, we generate the routing heatmaps, as shown in Fig. 7. It is evident that in the last two layers, i.e., $l=1$ and $l=2$, the gating network clusters data classes effectively, allowing each expert to specialize in handling specific classes. In contrast, the shallower layers do not exhibit clear expert specialization, suggesting that these MoE layers may not be necessary and that a single FFN can replace the role of multiple sparse experts. Based on this, we estimate the routing degree for CIFAR100 to be around 4 to 16. To validate this hypothesis, we experiment with the $L=2$ configuration, achieving an accuracy of 91.7%. This setup maintains good results while reducing parameters and improving efficiency.
@@
Table 5: Top-1 accuracy on CIFAR100 under different configurations.
$L=1$ $L=2$ $L=4$ $L=6$ $L=9$ $L=12$
w/o Shared Expert
$N=2$ 91.4 91.5 91.5 91.5 91.3 91.2
$N=4$ 91.4 91.5 91.3 90.7 89.2 78.4
$N=8$ 91.5 91.3 90.8 89.9 80.9 52.9
w/ Shared Expert
$N=2$ 91.5 91.6 91.7 91.7 91.6 91.6
$N=4$ 91.6 91.7 91.7 91.7 91.7 91.6
$N=8$ 91.6 91.6 91.7 91.7 91.7 91.5
@@
Layer Scanning. We further validate the results by layer scanning, as shown in Table 5. When no shared experts are employed, an unreasonable configuration of the number of MoE layers leads to significantly lower accuracy, which is even more pronounced than what we observed in ImageNet-1K. We attribute this to the fact that on datasets with smaller data volumes and fewer classes, overly sparse architectures hinder each expert from being sufficiently optimized. These results reinforce the necessity of incorporating shared experts to stabilize model convergence. Moreover, for the efficient ViMoE, the required routing degree (i.e., the number of expert combinations) is indeed smaller when the dataset contains fewer classes. It can be observed that incorporating MoE only in the deepest one or two layers is sufficient to achieve considerable accuracy.
Discussion. Comparing the CIFAR100 results with those from ImageNet-1K, we observe that fewer experts are required when there are fewer classes. This aligns with the intuition that having numerous experts handle simpler tasks does not provide additional benefits and may even introduce drawbacks. Therefore, training a smaller number of experts to be specialized and efficient is sufficient.
5 Related Work
Mixture-of-Experts (MoE) (Jacobs et al., 1991) has been widely studied for its ability to modularize learning and reduce interference across data domains (Lewis et al., 2021; Zhou et al., 2022; Rajbhandari et al., 2022; Gou et al., 2023; Xue et al., 2024; Zhu et al., 2024). MoE uses a gating network to assign which experts should handle each data sample. Early MoE models were densely activated, meaning every input triggered all experts, which, while functional, was computationally expensive due to the significant resources required to process each input through all experts (Masoudnia & Ebrahimpour, 2014). Modern mainstream MoE models can be regarded as an application of dynamic neural networks (Han et al., 2021), using sparse activation selecting only a subset of experts to handle each input, which greatly reduces computational costs while preserving model expressiveness and performance (Hwang et al., 2023; Hazimeh et al., 2021). This approach has become increasingly important in large language models, where efficiency and scalability are paramount. Notable works in NLP, such as Switch Transformers (Fedus et al., 2022), GShard (Lepikhin et al., 2020), and GLaM (Du et al., 2022), have successfully applied sparse MoE, demonstrating significant advancements in handling large-scale tasks while optimizing resource usage.
MoE in Vision Tasks. In recent years, the high efficiency of MoE in NLP tasks has motivated researchers to explore their applications in the visual domain. Works such as V-MoE (Riquelme et al., 2021) and M3vit (Fan et al., 2022) integrate sparse MoE architectures into Vision Transformers. By replacing certain dense feedforward layers with sparse MoE layers, these models achieve efficient modeling in image classification tasks, enhancing computational efficiency and performance. Simultaneously, pMoE (Chowdhury et al., 2023) and DiT-MoE (Fei et al., 2024) introduce sparse conditional computation mechanisms. Specifically, pMoE employs CNNs as experts, dynamically selecting image patches for each expert, thereby reducing computational costs while maintaining generalization performance. DiT-MoE optimizes input-dependent sparsity in large diffusion transformer models, improving the efficiency and performance of image generation. Additionally, AdaMV-MoE (Chen et al., 2023) and the work by (Wu et al., 2022) focus on multi-task visual recognition and efficient training of large MoE vision transformers.
Transformer for Vision. Transformer model initially achieved remarkable success in natural language processing and was later introduced into computer vision, leading to the development of Vision Transformers (ViT) (Dosovitskiy, 2020). ViT introduced a new approach to image processing by dividing images into patches and treating them like words in text, allowing for global feature extraction across the entire image. Unlike convolutional neural networks (CNNs) that rely on local receptive fields, ViT’s Transformer-based architecture captures broader context, achieving performance on par with, or exceeding, that of CNNs. In the realm of self-supervised learning, MoCov3 (Xinlei et al., 2021) extended the momentum contrastive learning approach to ViT, successfully training high-quality visual features from unlabeled data. Inspired by BERT’s (Kenton & Toutanova, 2019) masked language modeling, methods such as BEiT (Bao et al., 2021), MAE (He et al., 2022), and iBOT (Zhou et al., 2021) pre-train ViTs through masked image modeling to enhance the model’s generalization ability and representation learning. DINOv2 (Oquab et al., 2023) employed self-supervised learning methods based on knowledge distillation, utilizing larger datasets and longer training periods, allowing it to learn robust visual features in an unsupervised manner, further advancing self-supervised ViT.

### v2

@@ 4.2 Investigating Efficiency from Stable Structure
Routing Degree. Another interesting observation is that the number of MoE layers $L$ required varies with the number of experts $N$. We suggest this is related to the routing degree, which represents the number of possible expert combinations and can be simply defined as $D=(C^{k}_{N})^{L}$. Since we fix the gating selection to top-1 (i.e., $k=1$), we obtain $D=(C^{1}_{2})^{5}=32$ for $N=2$, $D=(C^{1}_{4})^{3}=64$ for $N=4$, and $D=(C^{1}_{8})^{2}=64$ for $N=8$. This implies that approximately $32$ to $64$ routing combinations are sufficient for effectively partitioning and processing the data. Fewer combinations may affect performance, while more do not yield further significant gains.
From another perspective, if we view the gating network allocating experts to data as a clustering process, the routing degree essentially reflects the number of clusters formed from the dataset. Each expert combination can then specialize in learning from the samples of its corresponding cluster, facilitating the model in reaching optimal effectiveness. Our results validate that end-to-end training can effectively achieve this clustering effect without the need for additional clustering strategies to provide prior information for the gating mechanism [29].
@@
Efficient ViMoE. The conclusions above are derived from scanning the number of MoE layers. From another perspective, we can approximate the routing degree by observing the expert allocations in each layer. As shown in Fig. 4, the routing heatmap provides evidence of which MoE layers play a critical role, potentially indicating the necessary expert combinations that impact the results. These insights guide us in refining the structural design, retaining the crucial MoE layers while removing the unnecessary ones, thereby developing a more efficient ViMoE.
@@
In Table 1, we present various ViMoE configurations and compare their parameter counts. Although sparse MoE layers increase the total number of parameters, since we set the gate to route each image to the top-1 expert, it achieves higher accuracy without increasing the activated parameter counts or the inference burden. With the inclusion of the shared expert, we further improve accuracy at a relatively low extra cost. For example, when $N=8$ and $L=2$, only 2.4M additional activated parameters are required to surpass the baseline by 1.1% in accuracy. Furthermore, a comparison with $L=12$ highlights the efficiency of our structural design for ViMoE, significantly reducing parameter count without sacrificing accuracy.
$N$ $L$ w/ Shared Expert Total Param. Activate Param. FLOPs Acc.
- 0 - 22.0M 22.0M 6.14G 83.1
2 5 27.9M 22.0M 6.14G 83.6
2 5 33.8M 27.9M 7.65G 84.3
2 12 50.4M 36.2M 9.77G 84.2
4 3 32.7M 22.0M 6.14G 83.9
4 3 36.2M 25.6M 7.05G 84.2
4 12 78.8M 36.2M 9.77G 84.2
8 2 38.6M 22.0M 6.14G 83.9
8 2 40.9M 24.4M 6.74G 84.2
8 12 135.5M 36.2M 9.77G 84.3
@@
$N$ w/ Shared Expert $L=1$ $L=2$ $L=3$ $L=6$ $L=12$
4 51.0 51.3 50.6 49.5 43.5
8 51.1 51.2 50.6 49.2 42.0
4 51.2 51.5 51.5 51.4 51.1
8 51.5 51.4 51.6 51.3 51.0
Table 2: Semantic segmentation (mIoU) on ADE20K under various configurations. The DINOv2 baseline gives 50.8 mIoU.
Figure 5: Routing heatmap of the $l$-th MoE layer for semantic segmentation on ADE20K, where $l=1$ represents the deepest (last) layer and $l=12$ denotes the shallowest (first) layer. Routing operates at the token level, where each image patch is allocated to an expert. The $x$-axis is the expert ID, and the $y$-axis is the class ID. The label order in each figure is adjusted for better readability. Darker colors indicate a higher proportion of images from the corresponding class routed to the expert.
5 Empirical Generalization of Observations
The above observations and conclusions are based on image classification. To demonstrate their generalizability, we conduct validation on semantic segmentation.
ViMoE Settings. When applying ViMoE to semantic segmentation, we adopt a routing approach at the token level (as described in Sec. 3.2), allowing different experts to specialize in distinct tokens, thereby achieving improved pixel-level classification results. For simplicity, a linear layer is trained to predict class logits from the patch tokens output by the last layer. It generates a low-resolution logit map (e.g., $37\times 37$ for a model with patch size $14$), which is then upsampled to the full resolution ($512\times 512$) to obtain a segmentation map [32]. More implementation details are provided in Sec. 6.2.
Baseline and Stable ViMoE. We use the DINOv2 [32] self-supervised pre-trained ViT-S/14 [9] and fine-tune it on ADE20K [50] as the baseline, which achieves 50.8 mIoU. As previously observed in image classification, ViMoE with shared experts tends to yield stable results, allowing for easier configuration of MoE layers. To verify whether this finding can be extrapolated to semantic segmentation, we chose the straightforward approach by applying MoE in every ViT block (i.e., $L=12$). Experiments are conducted with the number of experts set to $N=4$ and $N=8$, yielding 51.1 and 51.0 mIoU, respectively, as shown in the last column of Table 2. We also report results without shared experts for comparison, demonstrating that shared experts effectively mitigate the performance degradation associated with inappropriate expert configurations.
Routing Heatmap. We aim to observe the routing of tokens within the MoE layers to find evidence that multiple experts handle different pixel classes in a divide-and-conquer manner, similar to the approach we employ in image classification. Since each token in ViT corresponds to a $14\times 14$ patch rather than a single pixel, we partition the full resolution ($512\times 512$) segmentation label map into corresponding patches. Then, we assign the most frequently occurring label within each patch as the ground-truth class for the corresponding token. While this strategy may introduce inaccuracies at boundaries, the overall impact remains minimal. Based on this, we generate the routing heatmaps for ViMoE on the semantic segmentation task, as illustrated in Fig. 5, taking $N=8$ as an example. The routing patterns exhibit notable similarity to those in Fig. 4, validating the generalizability of our observations and conclusions from image classification. This evidence aligns with the expectation that multiple experts can specialize in processing different types of information.
Efficient Structures Derived from Observations. Based on the routing behavior across different layers shown in the heatmap, we can analyze the roles of individual experts. In deeper layers, the gating network effectively clusters data, allowing each expert to focus on specific classes. This observation indicates which layers in ViMoE play a critical role and which may be less essential. For the example with $N=8$, the final layer ($l=1$) exhibits strong expert specialization, whereas the shallower layers do not show this effect as prominently. Consequently, we experiment with using the MoE only in the final layer, replacing sparse experts in the remaining layers with the dense structure. The experimental results are presented in Table 2, where using $N=8$ and $L=1$ achieves performance advancing the baseline by 0.7 mIoU. Increasing the number of MoE layers does not yield further gains, which aligns with our previous conclusions. Moreover, since ADE20K contains 150 classes, the required number of expert combinations, i.e., routing degree, is lower compared to ImageNet-1K, which explains why fewer MoE layers can yield satisfactory results.
Discussion. When fewer classes exist, the required number of experts decreases accordingly, which is intuitively reasonable. Deploying many experts for more straightforward tasks provides no additional benefit and may even introduce drawbacks. Therefore, training a limited number of experts is sufficient to ensure specialization and efficiency.
Results Visualization. In Fig. 6, we present the semantic segmentation results of ViMoE (configured with $N=8$ and $L=1$) on ADE20K. Remarkably, the model achieves impressive results even with a linear layer as the mask decoder. Additionally, we map the expert allocation for each token in the MoE layer (i.e., $l=1$) back to the original image, where distinct colors represent different experts. This visualization highlights the specialization of experts and illustrates the task allocation mechanism when handling complex scenes. Specifically, each image patch is efficiently routed to the most appropriate expert, and objects with the same semantic class across different images are predominantly allocated to the same expert, echoing the conclusions drawn from Fig. 5.
Figure 6: Qualitative results of ViMoE for semantic segmentation on ADE20K. The expert allocation map shows that each image patch is effectively routed to the appropriate expert, and objects with the same semantic class across different images are predominantly allocated to the same expert. More results are shown in Fig. 9 and Fig. 10.
@@ 6.1 Image Classification on ImageNet-1K
Implementation Details. ViMoE is based on DINOv2 [32] and fine-tuned on ImageNet-1K [8] with $224\times 224$ image resolution. We train the small-size models for $200$ epochs with a peak learning rate of $1\times 10^{-4}$ and the base-size models for $100$ epochs with a peak learning rate of $5\times 10^{-5}$. We use the AdamW [38] optimizer with a batch size of $1024$, a weight decay of $0.05$, and a layer-wise learning rate decay of $0.65$. The MoE layer is configured with three numbers of experts ($N=2$, $N=4$, and $N=8$), selecting the top-1 expert, with the load balancing loss coefficient $\alpha$ set to $0.01$.
Method Arch. Activate Param. FLOPs Acc.
MoCov3 [45] ViT-S/16 22.1M 4.25G 81.4
DINO [49] ViT-S/16 22.1M 4.25G 81.5
BEiT [2] ViT-S/16 22.1M 4.25G 81.7
iBOT [51] ViT-S/16 22.1M 4.25G 82.3
DINOv2 [32] ViT-S/14 22.0M 6.14G 83.1
DINO [49] ViT-B/16 86.6M 17.58G 82.8
MoCov3 [45] ViT-B/16 86.6M 17.58G 83.2
MAE [19] ViT-B/16 86.6M 17.58G 83.6
BEiT [2] ViT-B/16 86.6M 17.58G 83.7
iBOT [51] ViT-B/16 86.6M 17.58G 84.4
DINOv2 [32] ViT-B/14 86.5M 23.19G 86.2
MoCov3 [45] ViT-L/16 304.3M 59.70G 84.1
MAE [19] ViT-L/16 304.3M 59.70G 85.9
BEiT [2] ViT-L/16 304.3M 59.70G 86.0
iBOT [51] ViT-L/16 304.3M 59.70G 86.6
ViMoE ViT-S/14 22.0M 6.14G 83.9
ViMoE ViT-S/14 24.4M 6.74G 84.2
ViMoE ViT-B/14 95.9M 25.61G 86.6
Table 3: Top-1 accuracy on ImageNet-1K. All models are evaluated at resolution $224\times 224$. We select $N=8$ and $L=2$ as a representative configuration for reporting.  indicates the inclusion of shared experts.
@@
Results. We compare ViMoE with various baseline methods based on the ViT architecture. As shown in Table 3, ViMoE achieves an 83.9% top-1 accuracy with ViT-S/14, which is 0.8% higher than DINOv2 without increasing activated parameters. With the inclusion of shared experts, the accuracy further improves to 84.2%, outperforming DINOv2 by 1.1%. Notably, the small-size ViMoE surpasses the performance of many base-size methods, and the base-size ViMoE achieves comparable results to other larger-size models, with less than one-third of the activated parameters. This is also illustrated in Fig. 1.
6.2 Semantic Segmentation on ADE20K
Implementation Details. We fine-tune ViMoE for $80$k iterations with a batch size of $32$ and a resolution of $512\times 512$ without using multi-scale training and testing. The learning rate is set to $5\times 10^{-5}$, and the load balancing loss coefficient $\alpha$ is set to $0.001$. We use a simple linear layer without an additional segmentation decoder. Other hyperparameters are kept consistent with those used in image classification.
Results. Table 4 demonstrates that ViMoE achieves performance superior to the DINOv2 baseline with only a slight increase in cost. Furthermore, by utilizing a simple linear-layer decoder, ViMoE significantly outperforms other methods, including those based on ViT-B/16, while requiring substantially less computational effort.
Method Arch. Decoder FLOPs mIoU
DeiT [39] ViT-S/16 UPerNet [44] 157G 44.5
iBOT [51] ViT-S/16 UPerNet [44] 157G 45.4
BEiT [2] ViT-B/16 UPerNet [44] 605G 45.8
DINO [49] ViT-B/16 UPerNet [44] 605G 46.8
MAE [19] ViT-B/16 UPerNet [44] 605G 48.1
iBOT [51] ViT-B/16 UPerNet [44] 605G 50.0
DINOv2 [32] ViT-S/14 Linear 47G 50.8
ViMoE ViT-S/14 Linear 50G 51.5
Table 4: Semantic segmentation on ADE20K. We select $N=8$ and $L=1$ with shared experts as a representative configuration for reporting. FLOPs metric is evaluated at resolution $512\times 512$.
Strategy $L$ $N$ Avg. # Experts Activate Param. Acc.
Token 2 8 14.3 $+$ 2Δ 38.9M 84.1
Token 3 4 11.4 $+$ 3Δ 35.5M 84.2
Token 5 2 9.8 $+$ 5Δ 33.6M 84.1
Token 12 8 93.6 $+$ 12Δ 132.6M 84.2
Image 2 8 2 $+$ 2Δ 24.4M 84.2
Image 3 4 3 $+$ 3Δ 25.6M 84.2
Image 5 2 5 $+$ 5Δ 27.9M 84.3
Table 5: Ablation studies of different routing strategies for image classification. The total number of experts is $(N+1)\times L$ (including one shared expert per layer). Δ denotes shared experts.
6.3 Ablation and Analysis
In this section, we conduct various ablation studies and analyses of ViMoE, primarily on image classification.
Routing Strategy. In Sec. 3.2, we propose aligning the routing strategy with the task objective, specifically selecting experts based on the entire image rather than individual tokens for image classification. In Table 5, we conduct an ablation study comparing these two strategies, showing no significant difference in accuracy. This indicates that the image-level routing strategy, while simpler, is effective as it aligns with the task objective of image classification. Additionally, the average number of routed experts and activated parameters per image confirms that image-level strategy is more efficient than token-level routing. For semantic segmentation, which requires pixel-level classification, an image-level MoE is evidently unsuitable. Therefore, we design only a token-level MoE to meet its requirements.
@@ 6.3 Ablation and Analysis
Comparison with Dense Structures. Previous results validate the advantage of the MoE structure over dense models. However, when we introduce the shared expert, activated parameters increase. To ensure fairness, we modify the DINOv2 baseline by aligning the number of activated parameters while maintaining a dense architecture. One feasible approach is to configure two experts in the MoE structure and select both, allowing an additional FFN to be incorporated within the ViT block. In Table 6, we compare dense structures with varying numbers of layers to sparse MoE configurations. While increasing the number of parameters yields accuracy gains, the sparse structure achieves superior performance with fewer activated parameters. For instance, the sparse MoE using only 24.4M activated parameters ($L=2$) outperforms the dense model with 36.2M activated parameters ($L=12$) by 0.3%.
Arch. $L$ $N$ Activate Param. FLOPs Acc.
Dense 0 - 22.0M 6.14G 83.1
Dense 2 - 24.4M 6.74G 83.6
Dense 3 - 25.6M 7.05G 83.8
Dense 5 - 27.9M 7.65G 83.8
Dense 12 - 36.2M 9.77G 83.9
Sparse 2 8 24.4M 6.74G 84.2
Sparse 3 4 25.6M 7.05G 84.2
Sparse 5 2 27.9M 7.65G 84.3
Table 6: Comparison between dense structure and sparse MoE. For dense structures, $L$ indicates that each of the last $L$ layers contains two FFNs to align the number of activated parameters.
@@
Routing Distribution. In Sec. 3.1, we introduce the load balancing loss to facilitate the training of sparse MoE models. It aims to ensure that multiple experts receive inputs more evenly, preventing degradation into a dense model due to most data being routed to a single expert. We calculate the proportion of data allocated to each expert in the MoE layers, as shown in Fig. 7, where the gating network distributes the data relatively evenly across multiple experts. Combined with the observations from Fig. 4, this validates the expectation that MoE layers enable different experts to handle specific information.
@@ 6.4 Validation on CIFAR100
In the previous Sec. 4, we derive insights and conclusions about image classification from experiments conducted on the ImageNet-1K [8] dataset. In this section, we further validate our ViMoE on the CIFAR100 [42] dataset.
@@
Implementation Details. The models are fine-tuned on CIFAR100 for $100$ epochs with a weight decay of $0.3$. The peak learning rate is set to $3\times 10^{-4}$ with a warm-up of $3$ epochs, while all other settings remain consistent with those used in the ImageNet-1K experiments.
$N$ w/ Shared Expert $L=1$ $L=2$ $L=4$ $L=6$ $L=9$ $L=12$
2 91.4 91.5 91.5 91.5 91.3 91.2
4 91.4 91.5 91.3 90.7 89.2 78.4
8 91.5 91.3 90.8 89.9 80.9 52.9
2 91.5 91.6 91.7 91.7 91.6 91.6
4 91.6 91.7 91.7 91.7 91.7 91.6
8 91.6 91.6 91.7 91.7 91.7 91.5
@@
Table 7: Top-1 accuracy on CIFAR100 under various configurations. The DINOv2 baseline gives a top-1 accuracy of 91.3%.
Figure 8: Routing heatmap of the $l$-th MoE layer for image classification on CIFAR100, where $l=1$ represents the deepest (last) layer and $l=12$ denotes the shallowest (first) layer. The $x$-axis is the expert ID, and the $y$-axis is the class ID. The label order in each figure is adjusted for better readability. Darker colors indicate a higher proportion of images from the corresponding class routed to the expert.
Baseline and Stable ViMoE. The DINOv2 [32] baseline with ViT-S/14 [9] achieves a top-1 accuracy of 91.3%. Considering that CIFAR-100 contains only 100 categories, a relatively small number of experts is sufficient, so we set $N=4$. Based on prior experience, ViMoE with the shared expert tends to yield stable results, allowing us more flexibility in setting the number of MoE layers. We opt for a straightforward configuration with $L=12$, and under this setup, ViMoE achieves a top-1 accuracy of 91.6%, surpassing the baseline by 0.3%. Additionally, we compare the model without shared experts, which yields an accuracy of only 78.4%, falling far short of the baseline. This demonstrates that MoE is not a simple design that guarantees stable gains. In fact, the optimization complexity introduced by sparse structures in certain ViT blocks may have significant negative impacts, further highlighting the necessity of designing a stable ViMoE.
Efficient Structures Derived from Observations. We observe the behavior of MoE within the stable ViMoE and further analyze which layers play a critical role. Following the approach outlined in Sec. 4.2, we generate the routing heatmaps, as shown in Fig. 8. It is evident that in the last two layers, i.e., $l=1$ and $l=2$, the gating network clusters data classes effectively, allowing each expert to specialize in handling specific classes. In contrast, the shallower layers do not exhibit explicit expert specialization, suggesting that these MoE layers may not be necessary and that a single FFN can replace the role of multiple sparse experts. Based on this, we estimate the routing degree for CIFAR100 to be around 4 to 16. To validate this hypothesis, we experiment with the $L=2$ configuration, achieving an accuracy of 91.7%. This setup maintains good results while reducing parameters and improving efficiency.
@@
Layer Scanning. We validate the ViMoE configuration through layer scanning, as shown in Table 7. When shared experts are not employed, inappropriate MoE layers lead to significantly lower accuracy, which is even more pronounced than what we observed in ImageNet-1K. We attribute this to the fact that on datasets with smaller data volumes and fewer classes, overly sparse architectures hinder each expert from being sufficiently optimized. These results reinforce the necessity of incorporating shared experts to stabilize model convergence. Moreover, for the efficient ViMoE, the required routing degree (i.e., the number of expert combinations) is indeed smaller when the dataset contains fewer classes. It can be observed that incorporating MoE only in the deepest one or two layers is sufficient to achieve considerable accuracy.